Even though I understand the foil process, we were asked a question like this;


I have the answer when solving for "a" which is 2, but how on earth did they get the answer. HELP! Can you I see the work on this type of problem. Thanks.

Answer by IA(6)

2a + 3b= 19
5a - 2b= 0
4a+6b=38 (double first equation)
15a-6b=0 (triple second equation)
to cancel out the b add the equations togther.
this gives 19a = 38. divide both sides by 19. a=2.
then substitute 2 into the equation for a and use the same method to solve b.